Circular No. 7637

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001cg IN IC 3900 W. D. Li, University of California at Berkeley, reports the discovery

by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 17.8) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on May 28.2 and 29.2 UT. The new object is located at α = 12^{h}55^{m}40^{s}.87, δ = +27^{o}15′03′′.9 (equinox 2000.0), which is 6′′.0 west and 1′′.5 north of the nucleus of IC 3900. A KAIT image taken on May 15.2 showed nothing at this position (limiting mag ∼ 19.0).

A. V. Filippenko and R. Chornock, University of California at Berkeley, write: “Inspection of a CCD spectrum (range 390–1000 nm), obtained on May 29 UT with the Keck-II 10-m reflector, reveals that SN 2001cg is a type-Ia supernova. The spectrum most closely resembles those of subluminous members of the class, like SN 1991bg and SN 1998de (e.g., Modjaz et al. 2001, PASP 113, 308), roughly 1–2 weeks past maximum brightness; the absorption at Si II 580-nm is unusually strong relative to the Si II 615nm line. The host galaxy (cz = 7100 km/s, from NED) appears to have a Hubble type of E or S0, consistent with the preponderance of early-type hosts among other subluminous type-Ia supernovae.”

SUPERNOVA 2001ch IN MCG –01-54-16 Li also reports another LOTOSS discovery of a supernova on unfiltered

KAIT images taken on May 28.5 (mag ∼ 16.7) and 29.5 UT (mag ∼ 16.8). SN 2001ch is located at α = 21^{h}25^{m}59^{s}.19, δ = −3^{o}48′46′′.8 (equinox 2000.0), which is ∼ 12′′ west and 14′′ south of the diffuse nucleus of MCG –01-54-16. A KAIT image taken on 2000 Nov. 28.3 showed nothing at this position (limiting mag ∼ 19.0).

Filippenko and Chornock add that a CCD spectrum (range 390–1000 nm), obtained on May 29 UT with the Keck-II 10-m reflector, reveals that SN 2001ch is a type-Ic supernova, just beginning to enter the nebular phase (perhaps 1–2 months past maximum brightness). The strength of [O I] 557.7-nm emission is comparable to that of [O I] 630.0-nm, but still weak compared with the Ca II near-infrared triplet. The O I 777.4-nm absorption line is strong, while lines of He I are weak or absent. The object is superposed on bright H II regions in the disk of the host galaxy, whose recession velocity is 2930 km/s (from NED).

2001 May 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green